An object of mass 1.67 kg is placed on a horizontal table and connected to a second object of mass 6.97 kg by a string passing over a pulley Determine the magnitude of the acceleration of the system, in m/s2. mj m2 Answer Check Determine the tension in the string, in newtons. Answer:
Answer #1

dynamics for hanging mass

m2 a = m2 g - T

T = m2 g - m2 a ... (i)

dynamics for other mass

T = m1 a

Putting in (i)

m1 a = m2 g - m2 a

a ( m1 + m2) = m2 g

a ( 1.67 + 6.97) = 6.97 * 9.8

a = 7.906 m/s^2

==========

puttinh in (i)

T = 13.202 N

=========

Comment in case any doubt, will reply for sure.. Goodluck